OpenStudy (anonymous):

A bicycle traveled 150 meters west from point A to point B. Then it took the same route and came back to point A. It took a total of 2 minutes for the bicycle to return to point A. What is the average speed and average velocity of the bicycle?

OpenStudy (shamim):

Average speed=distance / time=300/120=?

OpenStudy (shamim):

Average velocity=displacement/time=0/120=?

OpenStudy (shamim):

Can u tell me displacement = 0. Why?